OpenStudy (anonymous):

what is the value of c? 4x^2-32x+c. 8?

OpenStudy (anonymous):

because 8*4 is 32

OpenStudy (anonymous):

so that you have a perfect square?

OpenStudy (anonymous):

yes. I think so. a perfect square trinomial.

OpenStudy (anonymous):

\[4=2^2\] yes? and we want \[-32=-2\times 2\times c\]

OpenStudy (anonymous):

so yes 8 is right!

OpenStudy (anonymous):

(2x-8)^2.... c=64?

OpenStudy (anonymous):

Really? I got it right?

OpenStudy (anonymous):

would it be 8^2 for 64 or just 8?

OpenStudy (anonymous):

pmel is right it is 64

OpenStudy (nowhereman):

and I thought \[c \approx 200000 m/s^2\]

OpenStudy (anonymous):

b is 8 so b^2=64

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

pmel has it. you want a perfect square and it will look like \[(2x-b)^2\] \[b=8\] \[b^2=64\]

OpenStudy (anonymous):

in general, when working with perfect square quadratics the last term c is larger than the middle term b is ax^2+bx+c because for all numbers greater than 2 the value (arbitrary constant d) d^2>d+d.